# CompilerDream: Learning a Compiler World Model for General Code Optimization

### v3

@@ 1. Introduction
However, these learning-based methods still risk sacrificing some optimality and, more importantly, face a significant bottleneck in broad generalization across diverse programs that may be out of the training samples. A range of machine learning applications (Devlin et al., 2019; Brown et al., 2020; Radford et al., 2021; Kirillov et al., 2023) have witnessed that training high-capacity models on large-scale datasets can yield unprecedented performance. Prior studies on compiler optimization also suggest that training on large datasets with diverse programs could be beneficial (Cummins et al., 2022), yet the prevalent practice in this field is still to learn optimization strategies in a per-program manner (Haj-Ali et al., 2020; Shahzad et al., 2022) or from relatively small training sets comprising only a few hundred programs with limited-capacity models (Mammadli et al., 2020; Jain et al., 2022), which hinders generalization.
@@ 2. Related Work
Recent advancements have seen reinforcement learning (RL) techniques making strides in compiler optimization, circumventing the need for collecting optimal labeled data (Kulkarni and Cavazos, 2012). This technique has been applied to optimize individual compilation heuristics, such as inlining (Trofin et al., 2021), loop transformation (Haj-Ali et al., 2020; Brauckmann et al., 2024), and graph partitioning (Mirhoseini et al., 2017). Several works relevant to us, including AutoPhase (Haj-Ali et al., 2020), CORL (Mammadli et al., 2020), and POSET-RL (Jain et al., 2022), have explored the full optimization pipeline, i.e., the phase ordering problem, using model-free RL without incorporating a world model.
@@ 3.1. Phase Ordering Decision Process
Under the POMDP formulation, we design the observation and action space as summarized in Table 1. The observation features were selected for efficiency and effectiveness. In preliminary experiments, we find that complex program features like ProGraML (Cummins et al., 2021) and inst2vec (Ben-Nun et al., 2018) significantly slow down CompilerDream with marginal performance gains. In contrast, expert-designed Autophase features (Haj-Ali et al., 2020) leverage domain knowledge, enhancing generalization by filtering irrelevant details. We construct the observation by concatenating the 56-dimensional Autophase feature vector with a 42-dimensional action histogram vector, which records the number of times each action has been selected during the current episode. Both vectors are normalized for consistency: each Autophase feature is divided by the program’s initial total instruction count, and the action histogram is scaled by the per-episode action limit, which is set to 45. We adopt two distinct action spaces in our experiments to align with the baseline methods. The full action space consists of all 124 LLVM optimization passes, while the reduced action space derived from Autophase (Haj-Ali et al., 2020) consists of 42 actions. Originally, this action space included 45 LLVM passes, but CompilerGym excludes 3 due to updates in the latest LLVM version, leaving 42 actions. This reduced action space is widely used and has been shown to be effective in prior studies (Haj-Ali et al., 2020; Cummins et al., 2022).
@@
derived from Autophase(Haj-Ali et al., 2020).
@@ 3.3. Learning a Code Optimization Agent
The value of a program state $s$ is defined as the expected cumulative reward of applying an action sequence to the program. This aligns with the ratio of the reduced IR instruction count achieved between applying action sequence $\{a_{\tau}\}_{\tau=1}^{m}$ and standard $-Oz$ flag, according to the reward function $r$ defined in Eq. (1):
@@
[…]
@@
$(\displaystyle\mathcal{L}_{\text{actor}}(\psi)\doteq\mathbb{E}_{p_{\theta},\pi_{\psi}}\Big{[}\sum^{t+H}_{\tau=t}\Big{(}$$-\left(V_{\tau}-v_{\xi}(\hat{z}_{\tau})\right)\log\pi_{\psi}(\hat{a}_{\tau}\mid\hat{z}_{\tau})$
@@ 4.2. Autotuning: CompilerGym Leaderboard
60.8s 1.073$\times$
PPO + Guided Search 69.8s 1.070$\times$
CompilerDream 2.9s 1.068$\times$
Random Search ($t$=10800) 10,512.4s 1.062$\times$
Random Search ($t$=3600) 3,630.8s 1.061$\times$
Greedy Search 169.2s 1.055$\times$
GATv2 + DD-PPO 258.1s 1.047$\times$
@@ Results
As shown in Table 4, our RL agent trained on world model simulations achieves an average 1.068×1.068$\times$1.068 × code size reduction on cBench in a single trial, surpassing random search methods despite their longer execution times, even when accounting for CompilerDream’s training time. This demonstrates the world model’s ability to accurately simulate optimization processes and support effective agent learning. Furthermore, when combined with guided search, CompilerDream outperforms the top leaderboard method while using less wall time.
@@ 4.4. General Code Optimization with RL
(a)The performance of CompilerDream steadily and consistently improves throughout training, and achieves a validation code size reduction of 1.02 after about $4\times 10^{4}$ environment steps. In contrast, other methods (PPO, A2C, IMPALA, DQN, and APEX) all achieve reductions of less than 0.98 at the same stage. To reach a reduction of 1, PPO needs nearly $2\times 10^{5}$ steps–roughly six times the steps required by CompilerDream. The remaining four methods exhibit unstable performance and fail to exceed a reduction of 1 even after $10^{6}$ steps. (b) The generalization capabilities of different RL algorithms (CompilerDream, PPO, A2C, IMPALA, DQN, and APEX) on five test datasets unseen during training (cBench, CHStone, NPB, OpenCV, and TensorFlow). CompilerDream outperforms all other methods across all five datasets, demonstrating a clear advantage. PPO consistently ranks second, though with a noticeable performance gap compared to CompilerDream. The remaining four methods perform poorly, particularly on the cBench and CHStone datasets. (c) The impact of the training dataset is demonstrated by comparing CompilerDream trained on two datasets: CodeContests and Csmith. Across all five benchmarks (cBench, CHStone, NPB, OpenCV, and TensorFlow), CompilerDream trained on CodeContests consistently outperforms its counterpart trained on Csmith. Notably, the performance gap even reaches a geometric mean code size reduction difference of 0.2 on the CHStone benchmark.
@@
The performance of CompilerDream and PPO was evaluated on the large-scale FormAI test. For both methods, the code size reduction for most test cases (approximately $5\times 10^{4}$) falls within the range of $0.8\textasciitilde{}1.2$. As the reduction magnitude increases or decreases beyond this range, the number of corresponding cases gradually decreases. Comparing CompilerDream with PPO, we observe that CompilerDream optimizes more cases with higher code size reduction while producing fewer cases with worsened results. Notably, CompilerDream achieves a reduction of more than nine times for some cases, surpassing all previous studies.
@@
We train CompilerDream with its RL agent, and compare it with the following baselines: a random pass sequence, an autotuning approach using random search, LLVM’s -O0 and -Oz flag, and a state-of-the-art learning-based method (Haj-Ali et al., 2020) using Proximal Policy Optimization (PPO) (Schulman et al., 2017). The random search conducts hundreds of trials within a time budget similar to our RL agent. The -Oz flag represents LLVM’s highest level of code size optimization while the -O0 flag represents no optimization. The learning-based method using PPO and CompilerDream are trained on the large-scale CodeContests dataset and zero-shot generalized to unseen test programs. Following the state-of-the-art approach, all methods use the reduced action space (Table 1).
@@ Results
Moreover, CompilerDream’s zero-shot generalization matches or surpasses in-domain training. This advantage is particularly evident in the NPB dataset, where data sparsity limits in-domain agents, yet CompilerDream achieves an additional 3% code size reduction. Its robust generalization extends to new languages, as shown by results on the Fortran-based BLAS and NPB datasets. On the AI-generated Objective-C dataset, CompilerDream achieves an average code size reduction of 1.027$\times$, reaching up to 2.87$\times$ in certain test cases.
@@ Appendix A Extended Related Work
Machine learning techniques introduced in Section 2 require crafting high-quality features that capture the important characteristics of programs, a process known as feature engineering. The most prevalent feature vectors are based on the frequencies of various types of instructions or patterns within the programs (Magni et al., 2014; Haj-Ali et al., 2020), designed by expert intuitions. Numerous studies have aimed to reduce the cost of feature design. Following the success of word2vec embeddings in natural language processing (Mikolov et al., 2013), methods like code2vec (Alon et al., 2019), inst2vec (Ben-Nun et al., 2018), and IR2vec (VenkataKeerthy et al., 2020) represent programs as distributed vectors that capture syntactic and semantic information from the abstract syntax tree (AST) or intermediate representation (IR). The surge in deep learning has enabled feeding raw information such as AST (Alon et al., 2018), control-data flow graphs (CFG) (Brauckmann et al., 2020; Cummins et al., 2021), and code token sequences (Cummins et al., 2017) into powerful deep neural networks, capable of learning useful representations end-to-end.
@@ B.2. Details of Features and Actions
As shown in Table 1, we adpot the 56-dimensional Autophase feature and a reduced action space from Autophase (Haj-Ali et al., 2020). A detailed list of the Autophase features and actions space is available in Table 8 and Table 9.
@@ B.3. Hyperparameters
We use RLlib (Liang et al., 2018) to train and test model-free reinforcement learning algorithms including PPO (Schulman et al., 2017) (Haj-Ali et al., 2020), A2C (Mnih et al., 2016), IMPALA (Espeholt et al., 2018), APEX (Horgan et al., 2018), and DQN (Mnih et al., 2015). We use default hyperparameters of algorithms in RLlib following the CompilerGym platform (Cummins et al., 2022), except that we have carefully tuned the hyperparameters for PPO, as listed in Table 11. We explored roughly 2 or 3 values for each hyperparameter listed and selected the set of hyperparameters that yielded the best performance on the validation set. The results of the PPO baseline in Section 4.5 and Figure 6 are exactly the same as those in Section 4.4 and Figure 4.
@@ Appendix C Extended Experimental Results
cBench sha -sroa -gvn -sroa -sroa -instcombine -simplifycfg -sroa -functionattrs -licm -functionattrs -early-cse -simplifycfg -indvars -gvn -simplifycfg -early-cse -functionattrs -reassociate -memcpyopt -gvn -simplifycfg -reassociate -functionattrs -early-cse -instcombine -early-cse -lowerinvoke -simplifycfg -functionattrs -functionattrs -memcpyopt -gvn -simplifycfg -functionattrs -reassociate -early-cse -instcombine -early-cse -strip -strip -early-cse -strip -early-cse -strip -early-cse 799 349 500 (1.43$\times$)
bzip2 -sroa -gvn -sroa -sroa -simplifycfg -strip -strip -instcombine -simplifycfg -early-cse -simplifycfg -strip -strip -reassociate -reassociate -memcpyopt -jump-threading -memcpyopt -functionattrs -memcpyopt -functionattrs -licm -functionattrs -gvn -simplifycfg -functionattrs -functionattrs -reassociate -memcpyopt -early-cse -jump-threading -memcpyopt -early-cse -strip -strip -instcombine -simplifycfg -jump-threading -jump-threading -early-cse -jump-threading -early-cse -simplifycfg -strip -functionattrs 28748 13565 15946 (1.18$\times$)
OpenCV #41 -sroa -early-cse -early-cse -simplifycfg -early-cse -early-cse -early-cse -sink -sink -sink -loop-deletion -loop-deletion -loop-deletion -lowerinvoke -simplifycfg -sroa -sroa -sroa -early-cse -early-cse -early-cse -early-cse -lowerinvoke -lowerinvoke -lowerinvoke -globalopt -sroa -instcombine -simplifycfg -early-cse -early-cse -early-cse -sink -globalopt -sroa -sroa -functionattrs -simplifycfg -early-cse -functionattrs -early-cse -sink -functionattrs -functionattrs -functionattrs 28 18 28 (1.56$\times$)
#9 -sroa -gvn -sroa -early-cse -simplifycfg -strip -strip -loop-deletion -lowerinvoke -simplifycfg -sroa -early-cse -instcombine -simplifycfg -early-cse -memcpyopt -early-cse -strip -strip -strip -functionattrs -instcombine -simplifycfg -jump-threading -simplifycfg -memcpyopt -early-cse -simplifycfg -reassociate -memcpyopt -functionattrs -functionattrs -reassociate -memcpyopt -early-cse -instcombine -simplifycfg -jump-threading -early-cse -jump-threading -strip -simplifycfg -early-cse -simplifycfg -functionattrs 9510 6341 9269 (1.18$\times$)
TensorFlow #17 -sroa -gvn -early-cse -simplifycfg -early-cse -early-cse -instcombine -early-cse -sink -sink -sink -lowerinvoke -simplifycfg -sroa -early-cse -memcpyopt -early-cse -memcpyopt -instcombine -simplifycfg -early-cse -early-cse -early-cse -simplifycfg -early-cse -early-cse -strip -strip -indvars -simplifycfg -early-cse -memcpyopt -early-cse -instcombine -early-cse -early-cse -simplifycfg -early-cse -memcpyopt -early-cse -early-cse -instcombine -simplifycfg -early-cse -early-cse 5512 4247 5450 (1.28$\times$)
@@ C.1. Learning Curves
BLAS geomean 0.931 0.821$\pm$0.018 0.960$\pm$0.004 0.993$\pm$0.002 0.993$\pm$0.005 0.991$\pm$0.000
min 0.707 0.432$\pm$0.106 0.723$\pm$0.076 0.916$\pm$0.004 0.831$\pm$0.148 0.913$\pm$0.000
max 1.000 1.002$\pm$0.004 1.031$\pm$0.010 1.014$\pm$0.001 1.025$\pm$0.005 1.016$\pm$0.000
cBench geomean 0.481 0.737$\pm$0.015 0.858$\pm$0.020 1.010$\pm$0.004 N/A 1.036$\pm$0.002
min 0.299 0.401$\pm$0.097 0.560$\pm$0.030 0.817$\pm$0.109 N/A 0.962$\pm$0.001
max 0.626 1.150$\pm$0.113 1.298$\pm$0.083 1.405$\pm$0.014 N/A 1.395$\pm$0.054
CHStone geomean 0.487 0.701$\pm$0.043 1.037$\pm$0.006 1.071$\pm$0.002 N/A 1.094$\pm$0.005
min 0.402 0.278$\pm$0.108 0.889$\pm$0.044 0.967$\pm$0.027 N/A 1.000$\pm$0.006
max 0.655 1.038$\pm$0.033 1.304$\pm$0.038 1.313$\pm$0.045 N/A 1.378$\pm$0.021
Linux geomean 0.988 0.981$\pm$0.004 1.001$\pm$0.000 0.985$\pm$0.001 0.993$\pm$0.000 0.986$\pm$0.005
min 0.615 0.740$\pm$0.126 1.000$\pm$0.000 0.615$\pm$0.000 0.700$\pm$0.001 0.642$\pm$0.041
max 1.011 1.001$\pm$0.002 1.011$\pm$0.000 1.009$\pm$0.003 1.011$\pm$0.000 1.009$\pm$0.003
Mibench geomean 0.389 0.812$\pm$0.022 1.005$\pm$0.001 1.000$\pm$0.007 N/A 1.006$\pm$0.002
min 0.278 0.352$\pm$0.074 0.858$\pm$0.010 0.833$\pm$0.062 N/A 0.879$\pm$0.003
max 0.760 1.392$\pm$0.074 1.603$\pm$0.015 1.524$\pm$0.045 N/A 1.588$\pm$0.000
NPB geomean 0.530 0.833$\pm$0.011 1.074$\pm$0.010 1.071$\pm$0.018 1.075$\pm$0.017 1.108$\pm$0.001
min 0.191 0.414$\pm$0.043 0.805$\pm$0.104 0.801$\pm$0.070 0.810$\pm$0.082 0.886$\pm$0.000
max 1.066 1.848$\pm$0.193 2.141$\pm$0.062 2.235$\pm$0.101 2.315$\pm$0.028 2.343$\pm$0.012
OpenCV geomean 0.981 0.949$\pm$0.011 1.080$\pm$0.001 1.082$\pm$0.004 1.087$\pm$0.007 1.092$\pm$0.000
min 0.833 0.603$\pm$0.133 0.888$\pm$0.021 0.897$\pm$0.002 0.898$\pm$0.001 0.897$\pm$0.000
max 1.370 1.409$\pm$0.167 1.571$\pm$0.022 1.620$\pm$0.193 1.635$\pm$0.186 1.556$\pm$0.000
TensorFlow geomean 0.983 0.912$\pm$0.009 1.006$\pm$0.003 1.010$\pm$0.019 1.032$\pm$0.000 1.032$\pm$0.001
min 0.927 0.618$\pm$0.046 0.877$\pm$0.009 0.841±0plus-or-minus0$\pm$ 0± 0.134 0.970$\pm$0.010 0.968$\pm$0.011
max 1.010 1.235$\pm$0.034 1.267$\pm$0.007 1.281$\pm$0.006 1.289$\pm$0.008 1.282$\pm$0.002

### v4

@@ 1. Introduction
However, these learning-based methods still risk sacrificing some optimality and, more importantly, face a significant bottleneck in broad generalization across diverse programs that may be out of the training samples. A range of machine learning applications (Devlin et al., 2019; Brown et al., 2020; Radford et al., 2021; Kirillov et al., 2023) have witnessed that training high-capacity models on large-scale datasets can yield unprecedented performance. Prior studies on compiler optimization also suggest that training on large datasets with diverse programs could be beneficial (Cummins et al., 2022), yet the prevalent practice in this field is still to learn optimization strategies in a per-program manner (Haj-Ali et al., 2020b; Shahzad et al., 2022) or from relatively small training sets comprising only a few hundred programs with limited-capacity models (Mammadli et al., 2020; Jain et al., 2022), which hinders generalization.
@@ 2. Related Work
Recent advancements have seen reinforcement learning (RL) techniques making strides in compiler optimization, circumventing the need for collecting optimal labeled data (Kulkarni and Cavazos, 2012). This technique has been applied to optimize individual compilation heuristics, such as inlining (Trofin et al., 2021), loop transformation (Haj-Ali et al., 2020a; Brauckmann et al., 2024), and graph partitioning (Mirhoseini et al., 2017). Several works relevant to us, including AutoPhase (Haj-Ali et al., 2020b), CORL (Mammadli et al., 2020), and POSET-RL (Jain et al., 2022), have explored the full optimization pipeline, i.e., the phase ordering problem, using model-free RL without incorporating a world model.
@@ 3.1. Phase Ordering Decision Process
Under the POMDP formulation, we design the observation and action space as summarized in Table 1. The observation features were selected for efficiency and effectiveness. In preliminary experiments, we find that complex program features like ProGraML (Cummins et al., 2021) and inst2vec (Ben-Nun et al., 2018) significantly slow down CompilerDream with marginal performance gains. In contrast, expert-designed Autophase features (Haj-Ali et al., 2020b) leverage domain knowledge, enhancing generalization by filtering irrelevant details. We construct the observation by concatenating the 56-dimensional Autophase feature vector with a 42-dimensional action histogram vector, which records the number of times each action has been selected during the current episode. Both vectors are normalized for consistency: each Autophase feature is divided by the program’s initial total instruction count, and the action histogram is scaled by the per-episode action limit, which is set to 45. We adopt two distinct action spaces in our experiments to align with the baseline methods. The full action space consists of all 124 LLVM optimization passes, while the reduced action space derived from Autophase (Haj-Ali et al., 2020b) consists of 42 actions. Originally, this action space included 45 LLVM passes, but CompilerGym excludes 3 due to updates in the latest LLVM version, leaving 42 actions. This reduced action space is widely used and has been shown to be effective in prior studies (Haj-Ali et al., 2020b; Cummins et al., 2022).
@@
derived from Autophase(Haj-Ali et al., 2020b).
@@ 3.3. Learning a Code Optimization Agent
The value of a program state $s$ is defined as the expected cumulative reward of applying an action sequence to the program. This aligns with the ratio of the reduced IR instruction count achieved between applying action sequence $\{a_{\tau}\}_{\tau=1}^{m}$ and standard Oz-Oz flag, according to the reward function $r$ defined in Eq. (1):
@@
(8)  $\begin{split}&\text{ Actor: }\hat{a}_{t}\sim\pi_{\psi}\left(\hat{a}_{t}\mid\hat{z}_{t}\right)\\ &\text{ Critic: }v_{\xi}\left(\hat{z}_{t}\right)\approx\mathbb{E}_{p_{\theta},\pi_{\psi}}\left[\sum\nolimits_{\tau\geq t}\gamma^{\tau-t}\hat{r}_{\tau}\right].\end{split}$
@@
$\mathcal{L}_{\text{actor}}(\psi)\doteq\mathbb{E}_{p_{\theta},\pi_{\psi}}\Big[\sum^{t+H}_{\tau=t}\Big($ $-\left(V_{\tau}-v_{\xi}(\hat{z}_{\tau})\right)\log\pi_{\psi}(\hat{a}_{\tau}\mid\hat{z}_{\tau})$
@@ 4.2. Autotuning: CompilerGym Leaderboard
[…]
@@ Results
As shown in Table 4, our RL agent trained on world model simulations achieves an average $1.068\times$ code size reduction on cBench in a single trial, surpassing random search methods despite their longer execution times, even when accounting for CompilerDream’s training time. This demonstrates the world model’s ability to accurately simulate optimization processes and support effective agent learning. Furthermore, when combined with guided search, CompilerDream outperforms the top leaderboard method while using less wall time.
@@ 4.4. General Code Optimization with RL
We train CompilerDream with its RL agent, and compare it with the following baselines: a random pass sequence, an autotuning approach using random search, LLVM’s -O0 and -Oz flag, and a state-of-the-art learning-based method (Haj-Ali et al., 2020b) using Proximal Policy Optimization (PPO) (Schulman et al., 2017). The random search conducts hundreds of trials within a time budget similar to our RL agent. The -Oz flag represents LLVM’s highest level of code size optimization while the -O0 flag represents no optimization. The learning-based method using PPO and CompilerDream are trained on the large-scale CodeContests dataset and zero-shot generalized to unseen test programs. Following the state-of-the-art approach, all methods use the reduced action space (Table 1).
@@ Results
[…]
@@ Appendix A Extended Related Work
Machine learning techniques introduced in Section 2 require crafting high-quality features that capture the important characteristics of programs, a process known as feature engineering. The most prevalent feature vectors are based on the frequencies of various types of instructions or patterns within the programs (Magni et al., 2014; Haj-Ali et al., 2020b), designed by expert intuitions. Numerous studies have aimed to reduce the cost of feature design. Following the success of word2vec embeddings in natural language processing (Mikolov et al., 2013), methods like code2vec (Alon et al., 2019), inst2vec (Ben-Nun et al., 2018), and IR2vec (VenkataKeerthy et al., 2020) represent programs as distributed vectors that capture syntactic and semantic information from the abstract syntax tree (AST) or intermediate representation (IR). The surge in deep learning has enabled feeding raw information such as AST (Alon et al., 2018), control-data flow graphs (CFG) (Brauckmann et al., 2020; Cummins et al., 2021), and code token sequences (Cummins et al., 2017) into powerful deep neural networks, capable of learning useful representations end-to-end.
@@ B.2. Details of Features and Actions
As shown in Table 1, we adpot the 56-dimensional Autophase feature and a reduced action space from Autophase (Haj-Ali et al., 2020b). A detailed list of the Autophase features and actions space is available in Table 8 and Table 9.
@@ B.3. Hyperparameters
We use RLlib (Liang et al., 2018) to train and test model-free reinforcement learning algorithms including PPO (Schulman et al., 2017) (Haj-Ali et al., 2020b), A2C (Mnih et al., 2016), IMPALA (Espeholt et al., 2018), APEX (Horgan et al., 2018), and DQN (Mnih et al., 2015). We use default hyperparameters of algorithms in RLlib following the CompilerGym platform (Cummins et al., 2022), except that we have carefully tuned the hyperparameters for PPO, as listed in Table 11. We explored roughly 2 or 3 values for each hyperparameter listed and selected the set of hyperparameters that yielded the best performance on the validation set. The results of the PPO baseline in Section 4.5 and Figure 6 are exactly the same as those in Section 4.4 and Figure 4.
@@ Appendix C Extended Experimental Results
[…]
@@ C.1. Learning Curves
[…]